Help PLEASE!

5. Find M.M.E for the parameter 0 There are 3 observations, X1 = 0.1, X2 0.5, X3 = 0.8 2 2x 0x<e f(x) = 02
Answer #1

= ( xf(x)da = $*- - . de 1 cdr

By method of moment of estimation; 6 = 3X = (Xi) = X1 + X2 + X3 = 0.1 +0.5 +0.8 = 1.4